# Understanding Performance of Long-Document Ranking Models through Comprehensive Evaluation and Leaderboarding

### v2

@@ Abstract.
We evaluated 20+ Transformer models for ranking of long documents (including two recent LongP models trained with FlashAttention) and compared them with simple FirstP baselines, which consist in applying the same model to input truncated to the first 512 tokens (to satisfy the input-sequence constraint of most off-the-shelf Transformer models). We used MS MARCO Documents v1 as a primary training set and evaluated models in the zero-shot scenario as well as after fine-tuning on other collections.
@@
In our initial experiments with MS MARCO, TREC DLs, and Robust04 we found that long-document models underperformed FirstP or outperformed it only by a modest margin (at most 5% on average in terms of MRR or NDCG). We then conjectured that this was not due to models’ inability to process long context, but rather due to a positional bias of relevant passages, which tended to be among the first 512 document tokens. We found evidence that this bias was, indeed, present in at least two test sets, which motivated us to create a new collection MS MARCO FarRelevant (based on MS MARCO Passages) where the relevant passages were not present among the first 512 tokens.
@@
Unlike standard collections where we observed both little benefit from incorporating longer contexts and limited variability in model performance (within a few %), experiments on MS MARCO FarRelevant uncovered dramatic differences among models. To begin with, the FirstP models performed roughly at the random-baseline level in both zero-shot and fine-tuning scenarios. Second, simple aggregation models including MaxP and PARADE Attention had good zero-shot accuracy, but benefited little from fine-tuning. In contrast, most other models had poor zero-shot performance (sometimes at a random baseline level), but outstripped MaxP by as much as 13-28% after finetuning. We believe this indicates that the positional bias not only diminishes benefits of processing longer document contexts, but also leads to model overfitting to positional bias and performing poorly in a zero-shot setting when the distribution of relevant passages changes substantially.
We make our software and MS MARCO FarRelevant available at https://anonymous.4open.science/r/long_doc_rank_model_analysis_v2-78E9/.
Neural IR, long document ranking
ccs: Information systems Retrieval models and ranking
@@
Figure 1. Average relative gain (in %) vs. relative increase in run-time compared to respective FirstP baselines on MS MARCO, TREC DL 2019-2021, and Robust04 (for a representative subset of models).
1. Introduction
@@
Transformer models (Vaswani et al., 2017)—such as BERT (Devlin et al., 2019)—pretrained in a self-supervised manner considerably advanced state-of-the-art of core natural language processing (NLP) (Devlin et al., 2019; Radford et al., 2018) and information retrieval (Nogueira and Cho, 2019). However, due to quadratic cost of the self-attention with respect to an input sequence length, a number of “chunk-and-aggregate” approaches were proposed and evaluated (Dai and Callan, 2019; MacAvaney et al., 2019; Boytsov and Kolter, 2021; Li et al., 2024), but existing studies typically have at least one of the following shortcomings:
@@
  • Reliance only on small-scale query collections such as TREC DL (Craswell et al., 2020, 2022), Robust04 (Voorhees, 2004), and Gov2 Terabyte (Clark et al., 2005);
@@
  • Lacking comparison with LongP models—directly supporting long inputs—such as sparse-attention models Longformer and BigBird (Beltagy et al., 2020; Zaheer et al., 2020), or more recent full-attention models trained with FlashAttention (Dao et al., 2022);
@@
To fill this gap we evaluated over 20 recent models for ranking of long documents and carried out their systematic comparison using two popular document collections (MS MARCO Documents v1/v2 (Craswell et al., 2020) and Robust04 (Clarke et al., 2004)), diverse query sets (both large and small) and multiple training seeds. We found that ranking models capable of processing long documents—including LongP models with sparse or full attention—showed little to no improvement compared to their respective FirstP baselines (which truncated documents to satisfy the input-sequence constraint of most off-the-shelf Transformer models, i.e., 512 tokens).
This finding is generally in line with previously reported results (see § 3.3) and an ablation experiment showed that limited improvement over FirstP was not related to the choice of the backbone Transformer model (see Table 5). Furthermore, we used our best models to produce several high-ranking runs on a competitive leaderboard. This, in our view, strengthens the credibility of our evaluation.
From the efficiency-effectiveness plot in Fig. 1, we can see that all long-document models are at least 2$\times$ slower than respective FirstP baselines. The biggest average gain of merely 5% is achieved by the PARADE Attn model (with a BERT-base backbone) at the cost of being 2.5$\times$ slower than its FirstP variant. All LongP models are even slower and show less improvement. Given such small benefits at the cost of a substantial slow-down, one could question practicality of such models and suggest using FirstP variants instead.
@@
To further support the relevance-bias hypothesis and answer RQ3, we constructed a new synthetic collection MS MARCO FarRelevant where relevant passages were not present among the first 512 tokens. Using MS MARCO FarRelevant, we evaluated zero-shot transferred as well as fine-tuned models and found the following:
@@
  • Not only positional bias diminished benefits of processing longer document contexts, but also lead to models’ overfitting to this bias and performing poorly in a zero-shot setting when the distribution of relevant passages changed substantially (RQ3);
@@
  • We contributed to the nascent field of analytical experimentation with a full control of outcomes by creating a new dataset MS MARCO FarRelevant, which we made available together with code https://anonymous.4open.science/r/long_doc_rank_model_analysis_v2-78E9/.
  • Our study confirmed superiority of PARADE (Li et al., 2024) models, but also showed their limited benefits on standard collections, which we attributed to the existence of positional bias of relevant passages (in such collections);
@@
2. Methods
2.1. Related Work
@@
Neural Ranking models have been a popular topic in recent years (Guo et al., 2019), but the success of early approaches was controversial (Lin, 2019). This changed with an introduction of a bi-directional encoder-only Transformer model BERT (Devlin et al., 2019), which was a successor of GPT (Radford et al., 2018) and ELMO (Peters et al., 2018). BERT was hugely successful and its resounding success can be attributed to a combination of the large model size and massive pretraining using self-supervision. A number of different Transformer models such as ELECTRA (Clark et al., 2020), and DEBERTA (He et al., 2021) improve upon BERT using different training strategies and/or datasets. However, due to their architectural similarities we—following Lin et al (Lin et al., 2021)—collectively call them as BERT models.
@@
Nogueira and Cho were first to apply BERT to ranking of text documents (Nogueira and Cho, 2019). In the big-data regime—most notably in the TREC deep learning track (Craswell et al., 2020)—BERT-based models outperformed prior neural and non-neural approaches by a large margin. They were also quite successful for several small-scale query collections outperforming previous neural and traditional approaches (Li et al., 2024; MacAvaney et al., 2019; Dai and Callan, 2019).
@@
Despite their impressive performance, neural models are susceptible to distributions shift and learning superficial features. Several authors found that neural rankers applied to out-of-domain data do not outperform BM25 (Mokrii et al., 2021; Thakur et al., 2021). They can also be confused by superficial text modifications such as adding distractor sentences (MacAvaney et al., 2022). Likewise, ranking performance can decrease if a query is reformulated (Penha et al., 2022). Weller et al. (Weller et al., 2023) showed that neural models are not effective to “spot” negation and often perform at random level in this respect. However, we are not aware of the prior work studying robustness to positional biases of relevant passages.
@@
The Transformer model (Vaswani et al., 2017) uses an attention mechanism (Bahdanau et al., 2015) where each sequence position can attend to all the positions in the previous layer. Because self-attention complexity is quadratic with respect to a sequence length, direct processing of long documents is not always practical. Thus, a vast majority of existing Transformer models limit the input length to be at most 512 (subword) tokens.
@@
Until recently, there have been two general approaches to handling long documents: localization of attention and splitting documents into chunks each of which is processed separately. Attention-localization approaches combine a limited-span (i.e., a sliding window) attention with some form of a selective global attention. There are many such approaches proposed (see, e.g., a recent survey (Tay et al., 2020)) and it would be infeasible to evaluate them all. Instead we consider two popular models: Longformer (Beltagy et al., 2020) and Big-Bird (Zaheer et al., 2020).
@@
With a document-splitting approach, one has to split documents into several chunks, process each chunk separately, and aggregate results, e.g., by computing a maximum or a weighted prediction score (Yilmaz et al., 2019; Dai and Callan, 2019). With respect to training approaches, the MaxP and SumP models by Dai and Callan (Dai and Callan, 2019) assume that each chunk in a relevant document is relevant. However, this assumption is problematic as the degree of relevance varies from passage to passage. Yilmaz et al. (Yilmaz et al., 2019) work around this problem by training a MaxP BERT model on short documents and zero-transfer it to long documents. In this study we work around this problem by training all document-splitting approaches including MaxP (Dai and Callan, 2019) in the end-to-end fashion, i.e., by plugging aggregated document-level scores directly into a loss function (analogous to training of CEDR (MacAvaney et al., 2019) and PARADE (Li et al., 2024) models).
@@
More recently, it has also become possible to train longer-context models using FlashAttention (Dao et al., 2022). FlashAttention computes attention exactly and it does not eliminate quadratic complexity. However, it dramatically speeds ups training while reducing memory requirements by using an IO-efficient computation approach.
@@
Because our primary focus is accuracy and we aim to understand the limits of long-document models, we exclude from evaluation several recent models (e.g., (Hofstätter et al., 2021; Zou et al., 2021)) that achieve better efficiency-effectiveness trade-offs by pre-selecting certain document parts and feeding only selected parts into a BERT ranker.
2.2. Ranking with Long-Document Models
In this section, we describe long-document BERT models in more details. We assume that an input text is split into small chunks of texts called tokens. Although tokens can be complete English words, Transformer models usually split text into sub-word units (Wu et al., 2016).
The length of a document $d$—denoted as $|d|$—is measured in the number of tokens. Because neural networks cannot operate directly on text, a sequence of tokens $t_{1}t_{2}\ldots t_{n}$ is first converted to a sequences of $d$-dimensional embedding vectors $w_{1}w_{2}\ldots w_{n}$ by an embedding network. These embeddings are context-independent, i.e., each token is always mapped to the same vector (Collobert et al., 2011; Mikolov et al., 2013).
For a detailed description of Transformer models, please see the annotated Transformer guide (Rush, 2018) as well as the recent survey by Lin et al. (Lin, 2019), which focuses on the use of BERT-style cross-encoding models for ranking and retrieval. For this paper, it is necessary to know only the following basic facts:
  • BERT is an encoder-only model, which converts a sequence of tokens $t_{1}t_{2}\ldots t_{n}$ to a sequence of $d$-dimensional vectors $w_{1}w_{2}\ldots w_{n}$. These vectors—which are token representations from the last model layer—are commonly referred to as contextualized token embeddings (Peters et al., 2018);
  • BERT operates on word pieces (Wu et al., 2016) rather than on complete words;
  • Using a pooled representation of token vectors $w_{1}w_{2}\ldots w_{n}$, a linear layer is used to produce a ranking score. A nearly universal pooling approach in cross-encoding rankers is to use the vector of the [CLS]  token, i.e., $w_{1}$. However, we learned that some models (e.g., JINA (günther2023jina)) converge well only with mean pooling, i.e., they use ${1\over n}\sum_{i=1}^{n}w_{i}$.
A “vanilla” BERT ranker (dubbed as monoBERT by Lin et al. (Lin, 2019)) uses a single fully-connect layer $F$ as a prediction head, which converts the last-layer representation of the [CLS]  token (i.e., a contextualized embedding of [CLS]) into a scalar (Nogueira and Cho, 2019). It makes a prediction based on the following sequence of tokens: [CLS]  $q$ [SEP] $d$ [SEP], where $q$ is a query and $d$ is a document.
An alternative approach is to aggregate contextualized embeddings of regular tokens using a shallow neural network (MacAvaney et al., 2019; Boytsov and Kolter, 2021; Khattab and Zaharia, 2020) (possibly together with the contextualized embedding of [CLS]) . This was first proposed by MacAvaney et al. (MacAvaney et al., 2019) who also found that incorporating [CLS]  improves performance. However, Boytsov and Kolter proposed a shallow aggregating network that does not use the output of the [CLS]  token and achieved the same accuracy on MS MARCO datasets (Boytsov and Kolter, 2021).
Replacing the standard BERT model in the vanilla BERT ranker with a BERT variant that “natively” supports longer documents (e.g., Big-Bird (Zaheer et al., 2020)) is, perhaps, the simplest way to deal with long documents. We collectively call these models as LongP models. For a typical BERT model, however, long documents and queries need to be split or truncated so that the overall number of tokens does not exceed 512. In the FirstP mode, we process only the first chunk and ignore the truncated text. In the SplitP mode, each chunk is processed separately and the results are aggregated. In the remaining of this section, we discuss these approaches in detail.
2.2.1. LongP models
In our work, we benchmark both sparse-attention and full-attention models. Sparse attention LongP models include two popular options: Longformer (Beltagy et al., 2020) and Big-Bird (Zaheer et al., 2020). In that, we use the same approach to score documents as with the vanilla BERT ranker, namely, concatenating queries with documents and making a prediction based on the contextualized embedding of the [CLS]  token (Nogueira and Cho, 2019). Both Big-Bird and Longformer use a combination of the local, “scattered” (our terminology), and global attention. The local attention utilizes a sliding window of a constant length where each token attends to each other token within this window. In the case of the global attention, certain tokens can attend to all other tokens and vice-versa, In Big-Bird, only special tokens such as [CLS]  can attend globally. In Longformer, the user have to select such tokens explicitly. Following Beltagy et al. (Beltagy et al., 2020), who applied this technique to question-answering, we “place” global attention only on query tokens. Unlike the global attention, the scattered attention is limited to restricted sub-sets of tokens, but these subsets do not necessarily have locality. In Big-Bird the scattered attention relies on random tokens, whereas Longformer uses a dilated sliding-window attention with layer- and head-specific dilation.
Full-attention models include JINABert (günther2023jina) and MosaicBERT (Portes et al., 2023), henceforth, simply JINA and Mosaic. Both models use a recently proposed FlashAttention (Dao et al., 2022) to efficiently process long-contexts as well as special the positional embeddings AliBi (Press et al., 2022) that can generalize to document lengths not seen during training. In addition to minor architectural difference, the models differ in pretraining strategies. MOSAIC relies primarily on the masked language (MLM) objective without next sentence prediction (NSP). JINA uses this approach as a first step, following a RoBERTa pretraining strategy (Liu et al., 2019) and fine-tuning on retrieval and classification tasks with mean-pooled representations. We found that JINA lost an ability to effectively pool on the [CLS] token and we used mean-pooling instead. For MOSAIC we used pooling on [CLS].
2.2.2. SplitP models
Greedy partitioning into disjoint chunks
CEDR models (MacAvaney et al., 2019) and the Neural Model 1 (Boytsov and Kolter, 2021) use the first method, which involves:
  • tokenizing the document $d$;
MacAvaney et al. (MacAvaney et al., 2019) use contextualized embeddings as a direct replacement of context-free embeddings in the following neural architectures: KNRM (Xiong et al., 2017), PACRR (Hui et al., 2018), and DRMM (Guo et al., 2016). To boost performance, they incorporate [CLS]vectors in a model-specific way. We call the respective models as CEDR-KNRM, CEDR-PACRR, and CEDR-DRMM.
The Neural Model 1 (Boytsov and Kolter, 2021) uses the same greedy partitioning approach as CEDR, but a different aggregator network, which does not use the embeddings of the [CLS]  token. This network is a neural parametrization of the classic Model 1 (Berger and Lafferty, 1999; Brown et al., 1993).
Sliding window approach
The BERT MaxP/SumP (Dai and Callan, 2019) and PARADE (Li et al., 2024) models use a sliding window approach. Assume $w$ is the size of the window and $s$ is the stride. Then the processing can be summarized as follows:
  • tokenizing, the document $d$ into sub-words $t_{1}t_{2}\ldots t_{n}$ ;
BERT MaxP/SumP
These models (Dai and Callan, 2019) use a linear layer $F$ to produce $m$ relevance scores $F(cls_{i})$. Then complete document scores are computed as $\max_{i=1}^{m}F(cls_{i})$ and $\sum_{i=1}^{m}F(cls_{i})$ for the MaxP and SumP models, respectively.
PARADE
These models (Li et al., 2024) can be divided into two groups. The first group includes PARADE average, PARADE max, and PARADE attention, which all use simple approaches to produce an aggregated representation of $m$ [CLS]-vectors $cls_{i}$. To compute a relevance score these aggregated representations are passed through a linear layer $F$.
In particular, PARADE average and PARADE max combine $cls_{i}$ using averaging and the element-wise maximum operation, respectively to generate aggregated representation of $m$ [CLS]  tokens $cls_{i}$. [^1] Note that both PARADE average and AvgP vanilla ranker use the same approach to aggregate contextualized embeddings of [CLS] tokens, but they differ in their approach to select document chunks. In particular, AveP uses non-overlapping chunks while PARADE average relies on the sliding window approach. The PARADE attention model uses a learnable attention (Bahdanau et al., 2015) vector $C$ to compute a scalar weight $w_{i}$ of each $i$ as follows:
$w_{1}w_{2}\ldots w_{m}=softmax(C\cdot cls_{1},C\cdot cls_{2},\ldots,C\cdot cls_{m})$
These weights are used to compute the aggregated representation as $\sum_{i=1}^{m}w_{i}cls_{i}$
(1) $F(AggregTransf(C,cls_{1},cls_{2},\ldots,cls_{m})[0])$
Table 1. Ranking Performance on MS MARCO, TREC DL, and Robust04.
Retriever / Ranker MS MARCO TREC DL Robust04 Avg. gain
dev (2019-2021) title description over FirstP
MRR NDCG@10 NDCG@20
retriever 0.312 0.629 0.428 0.402
FirstP (BERT) 0.394 0.632 0.475 0.527
FirstP (Longformer) 0.404 0.643 0.483 0.540
FirstP (ELECTRA) 0.417 0.662 0.492 0.552
FirstP (DEBERTA) 0.415 0.672 0.534 0.596
FirstP (Big-Bird) 0.408 0.656 0.507 0.560
FirstP (JINA) 0.422 0.654 0.488 0.532
FirstP (MOSAIC) 0.423 0.643 0.453 0.538
PARADE Attn 0.416a $(+5.5\%)$ 0.652a $(+3.1\%)$ 0.503a $(+5.7\%)$ 0.556a $(+5.6\%)$ +$5.0\%$
Miscellaneous models
We attempted to implement additional state-of-the-art models (Gao and Callan, 2022; Fu et al., 2022). Gao and Callan (Gao and Callan, 2022) introduced a late-interaction model MORES+, which is a modular long document reranker that uses a sequence-to-sequence transformer in a non-auto-regressive mode. In MORES+ document chunks are first encoded using the encoder-only Transformer model. Then they use a modified decoder Transformer for joint query-to-all-document-chunk cross-attention: This modification changes a causal Transformer into an encoder-only bi-directional Transformer model. As of the moment of writing, the MORES+ model holds the first position on a competitive MS MARCO document leaderboard. [^2] https://microsoft.github.io/MSMARCO-Document-Ranking-Submissions/leaderboard/. However, the authors provide only incomplete implementation which does not fully match the description in the paper (i.e., crucial details are missing). We reimplemented this model to the best of our understanding, but our implementation failed to outperform even BM25.
Fu et al. (Fu et al., 2022) proposed a multi-view interactions-based ranking model (MIR). They implement inter-passage interactions via a multi-view attention mechanism, which enables information propagation at token, sentence, and passage levels. Due to the computational complexity, they restrict these interactions to a set of salient/pivot tokens. However, the paper does not provide enough details regarding the choices of these tokens. There is no software available and authors did not respond to our clarification requests. Thus, this model is also excluded from our evaluation.
We additionally implemented both the encoder-only and the encoder-decoder variant of LongT5 (Guo et al., 2022) as well as RoFormer (with ROPE embeddings) (Su et al., 2024). We eventually had to abandon them due to poor convergence (LongT5) and/or CUDA crashes (RoFormer).
3. Experiments
3.1. Data
Our primary datasets include two MS MARCO Documents collections (v1 and v2) (Bajaj et al., 2016; Craswell et al., 2020, 2022), Robust04 (Voorhees, 2004), and associated query sets. In addition, we created a collection MS MARCO FarRelevant by using passages and relevance judgments from the MS MARCO Passages collection. Document and query statistics are summarized in Tables 2 and 3. Robust04 is a small collection of 0.5M documents that has a mixture of news articles and government documents some of which are quite long. However it has only a small number of queries (250), which makes it a challenging benchmark for training models in a low-data regime. Each query has a title and a description, which represent a brief information need and a more elaborate request (often a proper English prose), respectively. We use Robust04 in a cross-validation settings with folds established by Huston and Croft (Huston and Croft, 2014) and provided via IR-datasets (MacAvaney et al., 2021).
Table 2. Document Statistics
data set # of documents avg. # of BERT tokens
@@ 3.1. Data
Table 3. Query Statistics

### v3

@@ Abstract
We evaluated 20+ Transformer models for ranking of long documents (including recent LongP models trained with FlashAttention) and compared them with a simple FirstP baseline, which applies the same model to the truncated input (at most 512 tokens). We used MS MARCO Documents v1 as a primary training set and evaluated both zero-shot transferred and fine-tuned models.
@@
On MS MARCO, TREC DLs, and Robust04 no long-document model outperformed FirstP by more than 5% in NDCG and MRR (when averaged over all test sets). We conjectured this was not due to models’ inability to process long context, but due to a positional bias of relevant passages, whose distribution was skewed towards the beginning of documents. We found direct evidence of this bias in some test sets, which motivated us to create MS MARCO FarRelevant (based on MS MARCO Passages) where the relevant passages were not present among the first 512 tokens.
@@
Unlike standard collections where we saw both little benefit from incorporating longer contexts and limited variability in model performance (within a few %), experiments on MS MARCO FarRelevant uncovered dramatic differences among models. The FirstP models performed roughly at the random-baseline level in both zero-shot and fine-tuning scenarios. Simple aggregation models including MaxP and PARADE Attention had good zero-shot accuracy, but benefited little from fine-tuning. Most other models had poor zero-shot performance (sometimes at a random baseline level), but outstripped MaxP by as much as 13-28% after finetuning. Thus, the positional bias not only diminishes benefits of processing longer document contexts, but also leads to model overfitting to positional bias and performing poorly in a zero-shot setting when the distribution of relevant passages changes substantially. We make our software and data available. [^1] https://anonymous.4open.science/r/long_doc_rank_model_analysis_v2-78E9/.
Understanding Performance of Long-Document Ranking Models through Comprehensive Evaluation and Leaderboarding
Leonid Boytsovthanks: Work done outside the scope of Amazon employment, primarily while employed by Bosch.thanks: Equal contribution. Amazon AWS AI Labs Pittsburgh, USA leo@boytsov.info                        David Akinpelu 22footnotemark: 2 Louisiana State University Baton Rouge, USA akinpeluakorede01@gmail.com
Tianyi Linthanks: Work done while studying at Carnegie Mellon University.and Fangwei Gao33footnotemark: 3and Yutian Zhao33footnotemark: 3and Jeffrey Huang33footnotemark: 3
Nipun Katyal  and Eric Nyberg Carnegie Mellon University Pittsburg, PA, USA
@@
Figure 1: Average relative gain (in %) vs. relative increase in run-time compared to respective FirstP baselines on MS MARCO, TREC DL 2019-2021, and Robust04 (for a representative subset of models).
Figure 2: Zero-shot vs. fine-tuned performance on MS MARCO FarRelevant for a representative set of models.
1 Introduction
@@
Transformer models Vaswani et al. (2017)—such as BERT Devlin et al. (2019)—pretrained in a self-supervised manner considerably advanced state-of-the-art of core natural language processing (NLP) Devlin et al. (2019); Radford et al. (2018) and information retrieval Nogueira and Cho (2019). However, due to quadratic cost of the self-attention with respect to an input sequence length, a number of “chunk-and-aggregate” approaches were proposed and evaluated Dai and Callan (2019); MacAvaney et al. (2019); Boytsov and Kolter (2021); Li et al. (2024), but existing studies typically have at least one of the following shortcomings:
@@
  • Reliance only on small-scale query collections such as TREC DL Craswell et al. (2020, 2022), Robust04 Voorhees (2004), and Gov2 Terabyte Clark et al. (2005);
@@
  • Lacking comparison with LongP models—directly supporting long inputs—such as sparse-attention models Longformer and BigBird Beltagy et al. (2020); Zaheer et al. (2020), or more recent full-attention models trained with FlashAttention Dao et al. (2022);
@@
To fill this gap we evaluated over 20 recent models for ranking of long documents and carried out their systematic comparison using two popular document collections: MS MARCO Documents v1/v2 Craswell et al. (2020) and Robust04 Clarke et al. (2004), diverse query sets (both large and small) and multiple training seeds. We found that ranking models capable of processing long documents—including LongP models with sparse or full attention—showed little to no improvement compared to their respective FirstP baselines (which truncated documents to satisfy the input-sequence constraint of most off-the-shelf Transformer models, i.e., 512 tokens).
This finding is generally in line with previously reported results (see § B.4) and an ablation experiment showed that limited improvement over FirstP was not related to the choice of the backbone Transformer model (see Table 7). Furthermore, we used our best models to produce several high-ranking runs on a competitive leaderboard. This, in our view, strengthens the credibility of our evaluation.
From the efficiency-effectiveness plot in Fig. 1, we can see that all long-document models are at least 2$\times$ slower than respective FirstP baselines. The biggest average gain of merely 5% is achieved by the PARADE Attn model (with a BERT-base backbone) at the cost of being 2.5$\times$ slower than its FirstP baseline. All LongP models are even slower and show less improvement. Given such small benefits at the cost of a substantial slow-down, one could question practicality of such models and suggest using FirstP variants instead.
@@
To further support the relevance-bias hypothesis and answer RQ3, we constructed a new synthetic collection MS MARCO FarRelevant where relevant passages were not present among the first 512 tokens. Using MS MARCO FarRelevant, we evaluated zero-shot transferred as well as fine-tuned models and found the following (see Fig. 2):
@@
  • In contrast, other long-document models had poor zero-shot performance (sometimes at a random baseline level), but outstripped respective MaxP baselines by as much as 13.3%-27.7% after finetuning (RQ3);
  • Not only positional bias diminished benefits of processing longer document contexts, but it also lead to models’ overfitting to the bias and performing poorly in a zero-shot setting when the distribution of relevant passages changed substantially (RQ3);
@@
  • We contributed to the nascent field of analytical experimentation with a full control of outcomes by creating a new dataset MS MARCO FarRelevant, which we made available together with code. [^2] https://anonymous.4open.science/r/long_doc_rank_model_analysis_v2-78E9/.
  • Our study confirmed superiority of PARADE Li et al. (2024) models, but also showed their limited benefits on standard collections, which we attributed to the existence of positional bias of relevant passages (in such collections);
@@
2 Methods
2.1 Related Work
@@
Neural Ranking models have been a popular topic in recent years Guo et al. (2019), but the success of early approaches was controversial Lin (2019). This changed with an introduction of a bi-directional encoder-only Transformer model BERT Devlin et al. (2019), which was a successor of GPT Radford et al. (2018) and ELMO Peters et al. (2018). BERT was hugely successful and its resounding success can be attributed to a combination of the large model size and massive pretraining using self-supervision. A number of different Transformer models such as ELECTRA Clark et al. (2020), and DEBERTA He et al. (2021) improve upon BERT using different training strategies and/or datasets. However, due to their architectural similarities we—following Lin et al Lin et al. (2021)—collectively call them as BERT models.
@@
Nogueira and Cho were first to apply BERT to ranking of text documents Nogueira and Cho (2019). In the big-data regime—most notably in the TREC deep learning track Craswell et al. (2020)—BERT models outperformed prior neural and non-neural approaches by a large margin. They were also quite successful for several small-scale query collections outperforming previous neural and traditional approaches Li et al. (2024); MacAvaney et al. (2019); Dai and Callan (2019).
@@
Despite their impressive performance, neural models are susceptible to the distribution shift and learning superficial features. Several authors found that neural rankers applied to out-of-domain data do not always outperform BM25 Thakur et al. (2021); Mokrii et al. (2021). They can also be confused by superficial text modifications such as adding distractor sentences MacAvaney et al. (2022). Likewise, ranking performance can decrease if a query is reformulated Penha et al. (2022). Weller et al. Weller et al. (2023) showed that neural models are not effective to “spot” negation and often perform at random level in this respect. However, we are not aware of the prior work systematically studying robustness to positional biases of relevant passages.
@@
The Transformer model Vaswani et al. (2017) uses an attention mechanism Bahdanau et al. (2015) where each sequence position can attend to all the positions in the previous layer. Because self-attention complexity is quadratic with respect to a sequence length, direct processing of long documents is not always practical. Thus, a vast majority of existing Transformer models limit the input length to be at most 512 (subword) tokens.
@@
Until recently, there have been two general approaches to handling long documents: localization of attention and splitting documents into chunks each of which is processed separately. Attention-localization approaches combine a limited-span (i.e., a sliding window) attention with some form of a selective global attention. There are many such approaches proposed (see, e.g., a survey by Tay et al. 2020) and it would be infeasible to evaluate them all. Instead we consider two popular models: Longformer Beltagy et al. (2020) and Big-Bird Zaheer et al. (2020).
@@
With a document-splitting approach, one has to split documents into several chunks, process each chunk separately, and aggregate results, e.g., by computing a maximum or a weighted prediction score Yilmaz et al. (2019); Dai and Callan (2019). With respect to training approaches, the MaxP and SumP models by Dai and Callan Dai and Callan (2019) assume that each chunk in a relevant document is relevant. However, this assumption is problematic as the degree of relevance varies from passage to passage. Yilmaz et al. Yilmaz et al. (2019) work around this problem by training a MaxP BERT model on short documents and zero-transfer it to long documents. In this study we work around this problem by training all document-splitting approaches including MaxP Dai and Callan (2019) in the end-to-end fashion, i.e., by plugging aggregated document-level scores directly into a loss function (analogous to training of CEDR MacAvaney et al. (2019) and PARADE Li et al. (2024) models).
@@
More recently, it has also become possible to train longer-context models using FlashAttention Dao et al. (2022). FlashAttention computes attention exactly and it does not eliminate quadratic complexity. However, it dramatically speeds ups training while reducing memory requirements by using an IO-efficient computation approach.
@@
Because our primary focus is accuracy and we aim to understand the limits of long-document models, we exclude from evaluation several recent models (e.g., Hofstätter et al. (2021); Zou et al. (2021)) that achieve better efficiency-effectiveness trade-offs by pre-selecting certain document parts and feeding only selected parts into a BERT ranker.
Recently, several teams have focused on creating challenging benchmarks for long-document retrieval. A recent LoCo v1 Saad-Falcon et al. (2024) benchmark has 12 datasets. Despite 11 out of 12 collections has average document lengths in the order of dozens of thousands tokens, the E5 model with a 512 token input limit achieves high NDCG@10 scores (in the range of 0.4-0.85) for seven out of 12 LoCo v1 datasets. This prompted Zhu et al., 2024 to propose a more challenging LongEmbed benchmark containing a mix of real and synthetic datasets Zhu et al. (2024).
Table 1: Distribution of Start/End Positions of Relevant Passages Inside Documents
input chunk #
MS MARCO dev
FIRA
Hofstätter et al. (2020b)
(crowd-sourced)
start end start end
1 85.9% 71.0% 83.8% 76.4%
2 9.1% 14.9% 9.9% 15.3%
3 2.6% 6.1% 2.3% 3.9%
4 1.2% 3.0% 2.2% 2.2%
5 0.6% 1.4% 0.7% 0.9%
6 0.6% 1.2% 0.4% 0.5%
6+ 0.1% 2.5% 0.7% 0.7%
  • Chunk size is 477 BERT tokens.
Table 2: Document Statistics
data set # of documents average # of
BERT tokens
@@
Table 3: Query Statistics